OpenStudy (anonymous):

another e integration: xe^(-x^2) x times e to the negative x squared

hartnn (hartnn):

substitute -x^2 = u then du =... ?

OpenStudy (jhannybean):

you have two functions, x and e^(something power) you're going to have to use integration by parts :) with u-sub :P I think!!

hartnn (hartnn):

if you observe, you'll find that integration by parts won't be required! :)

OpenStudy (jhannybean):

..I wasn't observing, lmao.

OpenStudy (anonymous):

is this the ans>e^-x^2

hartnn (hartnn):

you are missing some constants only.....

OpenStudy (anonymous):

u mean i shoud add a c as constant..........so the actual rslt wold be= e^-x^2

hartnn (hartnn):

no , no ...when you put u= -x^2 what did you get du =... ?

OpenStudy (anonymous):

oh sorry......i had e litl mistake................du=-2xdx.....so ans is=-1/2e^-x^2

hartnn (hartnn):

jhannybean is solving for you nicely, confirm your answer with her, and watch! LOL

OpenStudy (jhannybean):

\[\huge \int\limits xe^{(-x^2)}dx\]let u = -x^2 du = -2x dx -> du /-2 = xdx\[\huge -\frac12\int\limits e^{(u)}du = -\frac12e^{u}+ c \]\[\huge -\frac12e^{-x^{(2)}}+c\]

OpenStudy (jhannybean):

I hope you can read it. ^_^

hartnn (hartnn):

both of you are absolutely correct! :D

OpenStudy (anonymous):

thnx jhannybean......